You are considering investing in a​ zero-coupon bond that will pay you its face value of $1,000 in twelve years. If the bond is currently selling for $444.01​, then the internal rate of return​ (IRR) for investing in this bond is closest​ to:

A.6.0​%

B.8.1​%

C.9.2​%

D.7.0​%
Answer #1

IRR is calculated by using outflow = Inflow

444.01=1000/(1+r)12

  (1+r)12 = 1000/444.01

r = 7%